1-D Question 21 Solve the equation for exact solutions. arcsir(y - 3) } Posten . P67

2-Question 22 Use the fundamental identities to find the value of the trigonometric function. Find cose if tan 0-6 and sin @ <0
Answer #1

Sot (y-74) 273 Sm (pint lytalet) sm (173) y=74 B 2 y = 7+ 2 y = x+3/3 6 coso = 2 smoco, tanv=6 CasoLo , for tan oto 37 6 ox t